OpenStudy (anonymous):

Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can make corrections as many times as he wants. Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Phillip can receive. Identify this as converging or diverging.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

120% of 75, or 75×1.2 which is 90.

OpenStudy (anonymous):

that is what i got..is it correct?

OpenStudy (mathmale):

I'm thinking out loud here: The first term of this geometric series is 75. What is the common ratio, r? Is |r|<1, as it must be for convergence of this geometric series? We both need to think about how we'd find that common ratio, r.

OpenStudy (mathmale):

How would you express "two tenths of the missing points?" If the student's score is 75, how many points are missing, and how did you calculate this result?

OpenStudy (anonymous):

25 points are missing and .2?

OpenStudy (mathmale):

What would two tenths of the number of points missing / lost come to?

OpenStudy (anonymous):

5 which would give him 80...

OpenStudy (mathmale):

Yes, that does make sense: (2/10)(25)=50/10=5.

OpenStudy (mathmale):

Next: adding back those points, what's the current score?

OpenStudy (anonymous):

75+5=80?

OpenStudy (anonymous):

this is like the fifteenth time i have seen this exact same stupid problem where does it come from?

OpenStudy (mathmale):

Let's not label it "stupid," OK?

OpenStudy (anonymous):

i over think these kind of word problems

OpenStudy (anonymous):

you are going to add \[25\times \frac{1}{5}+25\times \left(\frac{1}{5}\right)^2+25\times \left(\frac{1}{5}\right)^3+...\]

OpenStudy (anonymous):

what is wrong with "stupid" ? looks stupid to me

OpenStudy (anonymous):

use \[\frac{a}{1-r}\] with \[a=5,r=\frac{1}{5}\]

OpenStudy (anonymous):

you can do it in your head what is \(1-\frac{1}{5}\) ?

OpenStudy (mathmale):

We started with 75 and added 5 points, which was 2/10 of the missing points, which was 2/10 of 25, or 5. @topstryker: See whether you can translate what we have done so far into satellite73's expression, which features 1/5 as the common ratio, r. It's important that you understand this (understand where that 1/5 came from). @satellite73: As helpers, I think we ought to shed as much POSITIVE thought on any given problem as is possible; it's our job to help (not hinder) the problem solver.

OpenStudy (anonymous):

@mathmale you are right, it is positively stupid still wondering where it comes from though

OpenStudy (mathmale):

Bad mood today, Sat? @topstryker: Satellite wants to know the CONTEXT in which you found this problem.

OpenStudy (anonymous):

would it be converging or diverging?..

OpenStudy (mathmale):

@topstryker : In your shoes my first question would be "where did that 1/5 come from?" After you're satisfied that you can answer that question, the next question would be "given the first term of this geometric series and the common ratio, r, how would I find the sum of this geometric series? There is an important criterion for convergence or divergence, and that's that the magnitude of the common ratio, r, must be less than 1.

OpenStudy (mathmale):

What is your "r" here? Is the magnitude of "r" less than 1, equal to 1 or greater than 1?

OpenStudy (anonymous):

1/5 have to do with 5?

OpenStudy (mathmale):

Where did we get 5 originally? Hint: 2/10 of the number of points missing. Mind going through that calculation again: number of points earned: 75. Number of points missing: ??? 2/10 of the numb er of points missing = ??

OpenStudy (mathmale):

This is nothing new. We've already done it: We started with 75 and added 5 points, which was 2/10 of the missing points, which was 2/10 of 25, or 5.

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

so you simplified the 2/10

OpenStudy (mathmale):

So we're going to add those 5 points to the original 75 to get ... how many points total, as the revised score? That looks logical...2/10 does result in 1/5.

OpenStudy (anonymous):

80 as the revised score

OpenStudy (mathmale):

Right. what's missing (how many points short of a perfect score are we)?

OpenStudy (anonymous):

20 now

OpenStudy (mathmale):

Right. And what's 2/10 of that?

OpenStudy (anonymous):

4

OpenStudy (mathmale):

Exactly. Add that to 80.

OpenStudy (anonymous):

84

OpenStudy (mathmale):

Right. So, 100-84, or 16, points are missing. What's 2/10 of that?

OpenStudy (mathmale):

Alternatively, what's 1/5 of 16?

OpenStudy (anonymous):

3.2

OpenStudy (mathmale):

right. If we're on the right track, we keep summing up all these increments in the score. If you were to use 75 as the first term, a, and 1/5 as the common ratio, r, and if the sum of the series is a ______ , then what is the sum of this series? This sum represents the max score 1-r that the student can earn if his original grade was 75.

OpenStudy (mathmale):

a=the first term of the series. r= the common ratio.

OpenStudy (anonymous):

93.75?

OpenStudy (anonymous):

r=.2?

OpenStudy (mathmale):

...which is the same as 1/5, right? Is the magnitude of 1/5 (the magnitude of 0.2) less than 1?

OpenStudy (anonymous):

yes it is less than 1

OpenStudy (mathmale):

If so, the series converges, and the sum is .... which is the highest grade the student can earn after a first grade of 75.

OpenStudy (mathmale):

So: think about what you have done. The hardest step, I thought, was identifying the value of r. Satellite73 got that right right away: r=1/5. a is the initial value, obviously 75.

OpenStudy (anonymous):

yes, only thing im confused is that formula you just gave me, answered 93.75..but i thought it was 80

OpenStudy (mathmale):

that 80 was the student's score after he/she had made corrections for the FIRST time. After the student had made corr. for the 2nd time, his/her score became 84, right?

OpenStudy (mathmale):

We go on and on until the poor student has made corrections a million times. (Just kidding).

OpenStudy (anonymous):

oh yes, because he can make as many changes?

OpenStudy (mathmale):

Right...he can try fixing up his test paper as many times as he wants.

OpenStudy (anonymous):

so the hightest score he can recieve is a 93.75?

OpenStudy (mathmale):

Yes, I believe that's the case. One last time: make certain you know how to identify the initial term of a geom. series and that you know how to figure / obtain the common ratio.

OpenStudy (mathmale):

If the common ratio has a magnitude of less than 1, the series converges, and the sum of this infinite series is a ------ 1 - r

OpenStudy (anonymous):

thank you so much! i really do appreciate it

OpenStudy (mathmale):

My great pleasure. Thanks for sticking with this as long as you did!